# Quasi-one-dimensional soliton in a self-repulsive spin-orbit-coupled dipolar spin-half and spin-one condensates

### v1

@@ I Introduction
The long-range non-local dipolar interaction allows the formation of different types of quasi-1D solitons in a BEC, which is not possible in the absence of a dipolar interaction. After the experimental observation of dipolar BECs of 52Cr 1 ; 2 ; 6 , 166Er 7 , 168Er 8 , and 164Dy 9 ; 11 atoms, with a large magnetic dipole moment, we find it highly relevant to perform a detailed study of quasi-1D solitons in an SO-coupled pseudo spin-half and spin-one dipolar BEC. There have been a few previous studies on different aspects of an SO-coupled dipolar BEC sodip1 ; sodip2 ; sodip3 ; sodip4 ; sodip5 ; sodip6 ; sodip7 ; sodip8 ; sodip9 ; yuka ; yuka1 as well as on the formation of a quasi-1D sg1 ; sm1 ; sm2 ; sm3 , quasi-two-dimensional (quasi-2D) quasi-2da ; quasi-2db and quasi-three-dimensional hpu soliton in an SO-coupled nondipolar BEC. These previous studies on SO-coupled nondipolar BEC solitons were performed on a self-attractive BEC with negative intraspecies and interspecies scattering lengths, where solitons are possible even in the absence of an SO coupling. The formation of symbiotic bright solitons has also been studied in self-repulsive binary BEC mixtures r7a ; r7b bound by interspecies attraction.
@@
The above-mentioned spatially-periodic modulation 14 ; na-solid ; baym1 ; baym2 ; 2020 ; stripe in density in a quasi-1D SO-coupled BEC soliton is a consequence of the formation of a supersolid s1 ; s2 ; s3 ; s4 ; s5 . A supersolid is a special quantum state of matter with a rigid spatially-periodic crystalline structure 54 , breaking continuous translational invariance, that flows with zero viscosity as a superfluid breaking continuous gauge invariance. Spatially-periodic supersolid stripe formation in a quasi-1D pseudo spin-half BEC of 23Na atoms has been experimentally observed na-solid .
@@ II.1 SO-coupled spin-half dipolar GP equation
where ${\mathbb{r}}\equiv\{x,y,z\}$, ${\boldsymbol{\rho}}\equiv\{y,z\}$, $\nabla_{\mathbb{r}}^{2}=(\partial^{2}/\partial x^{2}+\partial^{2}/\partial y^{2}+\partial^{2}/\partial z^{2})\equiv(\partial_{x}^{2}+\partial_{y}^{2}+\partial_{z}^{2})$, $p_{x}=-{\mbox{i}}\hbar\partial_{x}$ is the momentum along the $x$ axes, $\gamma$ is the strength of the SO-coupling interaction. We will consider the two possibilities of SO coupling given by $\eta=\sigma_{x}$ and $\eta=\sigma_{y}.$ Of these two types of SO couplings, the choice $\gamma\sigma_{y}p_{x}$ corresponds to an equal mixture of Dresselhaus dre and Rashba ras ; ras2 SO couplings and was realized in the pioneering experiments on SO coupling in a pseudo spin-half BEC of 87Rb exptso and 23Na exptso3 atoms. We will not consider the choice $\eta=\sigma_{z}$, as that choice does not permit any off-diagonal coupling and hence does not lead to the variety of solitons found with the choices (5) illustrated in Figs. 1-3 and leads to only bright-bright solitons (result not elaborated in this paper).
@@ III.1 SO-coupled spin-half GP equation
${\mbox{i}}\partial_{t}\psi_{j}=\frac{\delta E[\psi_{j}]}{\delta\psi_{j}}.$ (32)
@@ IV.1 Quasi-1D SO-coupled spin-half self-repulsive dipolar BEC soliton
To study the case of an attractive intraspecies interaction and a repulsive interspecies interaction we consider the parameters $c_{0}=-0.5,c_{2}=0.5,d=2.$ Although, the intraspecies interaction is attractive, in the absence of a dipolar interaction ($d=0$), and an SO coupling, there is no binary soliton formation in this case for the present set of parameters. However, there could be a separate (uncoupled) single-component soliton(s). For $\gamma=0.1$ we have the same dark-bright and bright-bright solitons, viz. Figs. 1(a) and 1(c), as illustrated in Figs. 2(a)-(b), in increasing order of energy. As $c_{0}\neq c_{2}$ in this case, the total densities and energies of the solitons in Figs. 2(a)-(b) are different. In the case of the dark-bright soliton of Fig. 2(a), the magnetization is almost unity (${|\cal M|}=0.97\approx 1$) indicating a very small number of atoms in the dark soliton component. However, the phase-separated bright-bright soliton of the type presented in Fig. 1(b) is not possible in this case for $\gamma=0.1$. For $\gamma=1$ we have (the same three types of solitons as presented in Fig. 1): the dark-bright, the phase-separated bright-bright and the overlapping bright-bright solitons, as illustrated in Figs. 2(c)-(e), respectively. Again, the dark-bright and phase-separated bright-bright solitons shown in Figs. 2(c)-(d) are quasi-degenerate ground states and have a spatially-periodic modulation in density, whereas the the overlapping bright-bright soliton of Fig. 2(e) is an excited state. The bright-bright solitons of Figs. 2(b) and 2(e) are possible only for SO coupling $\gamma\sigma_{x}p_{x}$ and not for $\gamma\sigma_{y}p_{x}$. In this case the component and total densities of the solitons in Figs. 2(b) and 2(e) for $\gamma_{1}=0.1$ and $\gamma_{2}=1$, respectively, are identical and the two energies $E_{1}=-0.041$ and $E_{2}=-0.536$ satisfy the relation (48). In Fig. 2(f) we plot the phase of the wave function of the two components of the dark-bright soliton presented in Fig. 2(a). As expected dsx , the phase of the dark component shows a jump of $\pi$ at the origin while that of the bright component is zero. In fact, there is such a jump whenever there is a zero in the wave function leading to many such jumps in the phase, for example, in the case of the soliton in Fig. 2(d).
@@ IV.2 Quasi-1D SO-coupled spin-one self-repulsive dipolar BEC soliton
We now consider the scenario of soliton formation in a quasi-1D self-repulsive SO-coupled dipolar spin-one BEC. In this case there are two contact interaction parameters $c_{0}$ and $c_{2}$. First, we consider an anti-ferromagnetic BEC originating from a positive $c_{2}>0$ ku . To study the case of an anti-ferromagnetic BEC, we consider the parameters $c_{0}=c_{2}=0.5,d=2.$ In this case also, for $d=\gamma=0$, there is no soliton formation. We find three types of solitons in this case: partially phase-separated bright-bright-bright, dark-bright-dark, and bright-dark-bright solitons as displayed in Figs. 4(a)-(c), in increasing order of energy. The partially phase-separated bright-bright-bright soliton can only be obtained with the SO coupling $\gamma\Sigma_{x}p_{x}$, all other solitons in this case can be obtained with both types of SO coupling $\gamma\Sigma_{x}p_{x}$ and $\gamma\Sigma_{y}p_{x}$. For a large $\gamma=1$, there are only following two quasi-degenerate solitons with a spatially-periodic modulation in density: dark-bright-dark and bright-dark-bright solitons illustrated in Figs. 4(d)-(e), respectively. In Fig. 4(f) we plot the phase of the wave function of the three components of the dark-bright-dark soliton presented in Fig. 4(a). As expected dsx the phase of the dark components shows a jump of $\pi$ at the origin while that of the bright component is zero.
@@
Next we study the formation of quasi-1D solitons in a ferromagnetic SO-coupled spin-one dipolar BEC employing the parameters $c_{0}=0.5,c_{2}=-0.1,d=2$. In this case also, for $d=\gamma=0$, there is no soliton formation. For $\gamma=0.1$, one can have only two types of distinct solitons, as illustrated in Fig. 5, (a) a bright-bright-bright soliton and (b) a phase-separated bright-bright-bright soliton. This phase-separated bright-bright-bright soliton was obtained by introducing a relative phase of $\pi$ between components $j=1$ and 3 in the initial state to generate a repulsion between these two components. Because of this additional repulsion, the phase-separated bright-bright-bright soliton of Fig. 5(b) appears as an excited state. The phase-separated soliton of Fig. 5(b) is distinct from the partially phase-separated bright-bright-bright soliton in the anti-ferromagnetic case, viz. Fig. 4(a). For $\gamma=0.1$, the quasi-1D solitons in the ferromagnetic BEC are completely distinct from the quasi-1D solitons in the anti-ferromagnetic case, viz. Fig. 4(a)-(c), with no dark soliton component in the ferromagnetic case. For a large SO-coupling strength ($\gamma=1$), the only possible soliton is of the bright-bright-bright type as shown in Figs. 5(c). In this case the component and total densities of the solitons in Figs. 5(a) and 5(e) for $\gamma_{1}=0.1$ and $\gamma_{2}=1$, respectively, are identical and the two energies $E_{1}=-0.026$ and $E_{2}=-0.521$ satisfy the relation (48).
@@ V Summary and Discussion
To search for a quasi-1D soliton in a uniform pseudo spin-half and spin-one SO-coupled dipolar, self-repulsive BEC, using the numerical solution of a mean-field model, we identify different types of solitons. In the absence of an SO coupling and a dipolar interaction, the net prevailing interaction in all the systems, considered in this paper, is repulsive and no solitons can be formed, which justifies the term self-repulsive. In the pseudo spin-half case, for repulsive intraspecies and interspecies contact interactions, for a small SO-coupling strength ($\gamma=0.1$), there are dark-bright, and phase-separated and overlapping bright-bright solitons, viz. Figs. 1(a)-(c). For an attractive intraspecies and a repulsive interspecies contact interaction, for a small $\gamma$, we have only dark-bright and bright-bright solitons, viz. Figs. 2(a)-(b). For a repulsive intraspecies and an attractive interspecies contact interaction, for all $\gamma$ (=0.1,=0.1, and $1$), only the overlapping bright-bright soliton of Fig. 1(c) is possible, viz. Fig. 3(a)-(b). For a large SO-coupling strength ($\gamma=1$), for repulsive interspecies interaction and both attractive and repulsive intraspecies interactions, the scenario of soliton formation is the same and we can have (A) a dark-bright soliton, (B) a phase-separated bright-bright soliton, and (C) an overlapping bright-bright soliton, viz. Figs. 1(d)-(f) and Figs. 2(c)-(e). The first two of these solitons, e.g. (A) and (B), are quasi-degenerate and also have a spatially-periodic modulation in density. The overlapping bright-bright soliton of Fig. 1(f) is also degenerate with the other two solutions, whereas that of Fig. 2(e) is an excited state.

### v3

@@ I Introduction
The long-range non-local dipolar interaction allows the formation of different types of quasi-1D solitons in a BEC, which is not possible in the absence of a dipolar interaction. After the experimental observation of dipolar BECs of 52Cr 1 ; 2 ; 6 , 166Er 7 , 168Er 8 , and 164Dy 9 ; 11 atoms, with a large magnetic dipole moment, we find it highly relevant to perform a detailed study of quasi-1D solitons in an SO-coupled pseudo spin-half and spin-one dipolar BEC. There have been a few previous studies on different aspects of an SO-coupled dipolar BEC sodip1 ; sodip2 ; sodip3 ; sodip4 ; sodip5 ; sodip6 ; sodip7 ; sodip8 ; sodip9 ; yuka ; yuka1 as well as on the formation of a quasi-1D sg1 ; sm1 ; quasi-1d ; sm3 , quasi-two-dimensional (quasi-2D) quasi-2da ; quasi-2db and quasi-three-dimensional hpu soliton in an SO-coupled nondipolar BEC. These previous studies on SO-coupled nondipolar BEC solitons were performed on a self-attractive BEC with negative intraspecies and interspecies scattering lengths, where solitons are possible even in the absence of an SO coupling. The formation of symbiotic bright solitons has also been studied in self-repulsive binary BEC mixtures r7a ; r7b bound by interspecies attraction.
@@
The above-mentioned spatially-periodic modulation 14 ; exptso3 ; baym1 ; baym2 ; 2020 ; stripe in density in a quasi-1D SO-coupled BEC soliton is a consequence of the formation of a supersolid s1 ; s2 ; s3 ; s4 ; s5 . A supersolid is a special quantum state of matter with a rigid spatially-periodic crystalline structure 54 , breaking continuous translational invariance, that flows with zero viscosity as a superfluid breaking continuous gauge invariance. Spatially-periodic supersolid stripe formation in a quasi-1D pseudo spin-half BEC of 23Na atoms has been experimentally observed exptso3 .
@@ II.1 SO-coupled spin-half dipolar GP equation
where ${\mathbb{r}}\equiv\{x,y,z\}$, ${\boldsymbol{\rho}}\equiv\{y,z\}$, $\nabla_{\mathbb{r}}^{2}=(\partial^{2}/\partial x^{2}+\partial^{2}/\partial y^{2}+\partial^{2}/\partial z^{2})\equiv(\partial_{x}^{2}+\partial_{y}^{2}+\partial_{z}^{2})$, $p_{x}=-{\mbox{i}}\hbar\partial_{x}$ is the momentum along the $x$ axes, $\gamma$ is the strength of the SO-coupling interaction. We will consider the two possibilities of SO coupling given by $\eta=\sigma_{x}$ and $\eta=\sigma_{y}.$ Of these two types of SO couplings, the choice $\gamma\sigma_{y}p_{x}$ corresponds to an equal mixture of Dresselhaus dre and Rashba ras SO couplings and was realized in the pioneering experiments on SO coupling in a pseudo spin-half BEC of 87Rb exptso and 23Na exptso3 atoms. We will not consider the choice $\eta=\sigma_{z}$, as that choice does not permit any off-diagonal coupling and hence does not lead to the variety of solitons found with the choices (5) illustrated in Figs. 1-3 and leads to only bright-bright solitons (result not elaborated in this paper).
@@ III.1 SO-coupled spin-half GP equation
${\mbox{i}}\partial_{t}\psi_{j}=\frac{\delta E[\psi_{j}]}{\delta\psi_{j}^{*}}.$ (32)
@@ IV.1 Quasi-1D SO-coupled spin-half self-repulsive dipolar BEC soliton
To study the case of an attractive intraspecies interaction and a repulsive interspecies interaction we consider the parameters $c_{0}=-0.5,c_{2}=0.5,d=2.$ Although, the intraspecies interaction is attractive, in the absence of a dipolar interaction ($d=0$), and an SO coupling, there is no binary soliton formation in this case for the present set of parameters. However, there could be a separate (uncoupled) single-component soliton(s). For $\gamma=0.1$ we have the same dark-bright and bright-bright solitons, viz. Figs. 1(a) and 1(c), as illustrated in Figs. 2(a)-(b), in increasing order of energy. The soliton in Fig. 2(b) is in an excited metastable state. As $c_{0}\neq c_{2}$ in this case, the total densities and energies of the solitons in Figs. 2(a)-(b) are different. In the case of the dark-bright soliton of Fig. 2(a), the magnetization is almost unity (${|\cal M|}=0.97\approx 1$) indicating a very small number of atoms in the dark soliton component. However, the phase-separated bright-bright soliton of the type presented in Fig. 1(b) is not possible in this case for $\gamma=0.1$. For $\gamma=1$ we have (the same three types of solitons as presented in Fig. 1): the dark-bright, the phase-separated bright-bright and the overlapping bright-bright solitons, as illustrated in Figs. 2(c)-(e), respectively. Again, the dark-bright and phase-separated bright-bright solitons shown in Figs. 2(c)-(d) are quasi-degenerate ground states and have a spatially-periodic modulation in density, whereas the the overlapping bright-bright soliton of Fig. 2(e) is an excited metastable state. The bright-bright solitons of Figs. 2(b) and 2(e) are possible only for SO coupling $\gamma\sigma_{x}p_{x}$ and not for $\gamma\sigma_{y}p_{x}$. In this case the component and total densities of the solitons in Figs. 2(b) and 2(e) for $\gamma_{1}=0.1$ and $\gamma_{2}=1$, respectively, are identical and the two energies $E_{1}=-0.041$ and $E_{2}=-0.536$ satisfy the relation (48). In Fig. 2(f) we plot the phase of the wave function of the two components of the dark-bright soliton presented in Fig. 2(a). As expected dsx , the phase of the dark component shows a jump of $\pi$ at the origin while that of the bright component is zero. In fact, there is such a jump whenever there is a zero in the wave function leading to many such jumps in the phase, for example, in the case of the soliton in Fig. 2(d).
@@ IV.2 Quasi-1D SO-coupled spin-one self-repulsive dipolar BEC soliton
We now consider the scenario of soliton formation in a quasi-1D self-repulsive SO-coupled dipolar spin-one BEC. In this case there are two contact interaction parameters $c_{0}$ and $c_{2}$. First, we consider an anti-ferromagnetic BEC originating from a positive $c_{2}>0$ ku . To study the case of an anti-ferromagnetic BEC, we consider the parameters $c_{0}=c_{2}=0.5,d=2.$ In this case also, for $d=\gamma=0$, there is no soliton formation. We find three types of solitons in this case: partially phase-separated bright-bright-bright, dark-bright-dark, and bright-dark-bright solitons as displayed in Figs. 4(a)-(c), in increasing order of energy. The solitons in Figs. 4(b)-(c) are excited metastable states. The partially phase-separated bright-bright-bright soliton can only be obtained with the SO coupling $\gamma\Sigma_{x}p_{x}$, all other solitons in this case can be obtained with both types of SO coupling $\gamma\Sigma_{x}p_{x}$ and $\gamma\Sigma_{y}p_{x}$. For a large $\gamma=1$, there are only following two quasi-degenerate solitons with a spatially-periodic modulation in density: dark-bright-dark and bright-dark-bright solitons illustrated in Figs. 4(d)-(e), respectively. In Fig. 4(f) we plot the phase of the wave function of the three components of the dark-bright-dark soliton presented in Fig. 4(a). As expected dsx the phase of the dark components shows a jump of $\pi$ at the origin while that of the bright component is zero.
@@
Next we study the formation of quasi-1D solitons in a ferromagnetic SO-coupled spin-one dipolar BEC employing the parameters $c_{0}=0.5,c_{2}=-0.1,d=2$. In this case also, for $d=\gamma=0$, there is no soliton formation. For $\gamma=0.1$, one can have only two types of distinct solitons, as illustrated in Fig. 5, (a) a bright-bright-bright soliton and (b) a phase-separated bright-bright-bright soliton. This phase-separated bright-bright-bright soliton was obtained by introducing a relative phase of $\pi$ between components $j=1$ and 3 in the initial state to generate a repulsion between these two components. Because of this additional repulsion, the phase-separated bright-bright-bright soliton of Fig. 5(b) appears as an excited metastable state. The phase-separated soliton of Fig. 5(b) is distinct from the partially phase-separated bright-bright-bright soliton in the anti-ferromagnetic case, viz. Fig. 4(a). For $\gamma=0.1$, the quasi-1D solitons in the ferromagnetic BEC are completely distinct from the quasi-1D solitons in the anti-ferromagnetic case, viz. Fig. 4(a)-(c), with no dark soliton component in the ferromagnetic case. For a large SO-coupling strength ($\gamma=1$), the only possible soliton is of the bright-bright-bright type as shown in Figs. 5(c). In this case the component and total densities of the solitons in Figs. 5(a) and 5(e) for $\gamma_{1}=0.1$ and $\gamma_{2}=1$, respectively, are identical and the two energies $E_{1}=-0.026$ and $E_{2}=-0.521$ satisfy the relation (48).
@@ V Summary and Discussion
To search for a quasi-1D soliton in a uniform pseudo spin-half and spin-one SO-coupled dipolar, self-repulsive BEC, using the numerical solution of a mean-field model, we identify different types of solitons. In the absence of an SO coupling and a dipolar interaction, the net prevailing interaction in all the systems, considered in this paper, is repulsive and no solitons can be formed, which justifies the term self-repulsive. In the pseudo spin-half case, for repulsive intraspecies and interspecies contact interactions, for a small SO-coupling strength ($\gamma=0.1$), there are dark-bright, and phase-separated and overlapping bright-bright solitons, viz. Figs. 1(a)-(c). For an attractive intraspecies and a repulsive interspecies contact interaction, for a small $\gamma$, we have only dark-bright and bright-bright solitons, viz. Figs. 2(a)-(b). For a repulsive intraspecies and an attractive interspecies contact interaction, for all $\gamma$ (=0.1,=0.1, and $1$), only the overlapping bright-bright soliton of Fig. 1(c) is possible, viz. Fig. 3(a)-(b). For a large SO-coupling strength ($\gamma=1$), for repulsive interspecies interaction and both attractive and repulsive intraspecies interactions, the scenario of soliton formation is the same and we can have (A) a dark-bright soliton, (B) a phase-separated bright-bright soliton, and (C) an overlapping bright-bright soliton, viz. Figs. 1(d)-(f) and Figs. 2(c)-(e). The first two of these solitons, e.g. (A) and (B), are quasi-degenerate and also have a spatially-periodic modulation in density. The overlapping bright-bright soliton of Fig. 1(f) is also degenerate with the other two solutions, whereas that of Fig. 2(e) is an excited metastable state.
@@
Some of the solitons, viz. Figs. 2(b), 2(e), 4(b)-(c), and 5(b), presented in this paper are excited metastable states trapped in a local minimum of energy. Their existence, stability, and eventual decay are governed by an interplay between energetic (thermodynamic) considerations and the evolution of perturbations (dynamical stability). The imaginary-time propagation procedure has been adapted successfully for these metastable states, starting with a trial wave function with a definite symmetry property. In addition, these solitons are dynamically stable, as verified by the real-time propagation procedure. The stability of these metastable states in both real- and imaginary-time procedures generally may assure the thermodynamic stability of these states, but this is not guaranteed. Being excited states, they could be thermodynamically unstable and suffer energetic decay to the ground state, with a global minimum of energy, and phonons. A full analysis of the thermodynamic stability of these states is beyond the scope of this paper.